# Robust Finetuning of Vision-Language-Action Robot Policies via Parameter Merging

### v1

@@ Abstract
Generalist robot policies, trained on large and diverse datasets, have demonstrated the ability to generalize across a wide spectrum of behaviors, enabling a single policy to act in varied real-world environments. However, they still fall short on new tasks not covered in the training data. When finetuned on limited demonstrations of a new task, these policies often overfit to the specific demonstrations—not only losing their prior abilities to solve a wide variety of generalist tasks but also failing to generalize within the new task itself. In this work, we aim to develop a method that preserves the generalization capabilities of the generalist policy during finetuning, allowing a single policy to robustly incorporate a new skill into its repertoire. Our goal is a single policy that both learns to generalize to variations of the new task and retains the broad competencies gained from pretraining. We show that this can be achieved through a simple yet effective strategy: interpolating the weights of a finetuned model with that of the pretrained model. We show, across extensive simulated and real-world experiments, that such model merging produces a single model that inherits the generalist abilities of the base model and learns to solve the new task robustly, outperforming both the pretrained and finetuned model on out-of-distribution variations of the new task. Moreover, we show that model merging enables continual acquisition of new skills in a lifelong learning setting, without sacrificing previously learned generalist abilities.
@@ 1 Introduction
Generalist robot policies trained on large corpora of data have recently shown impressive generalization abilities: out of the box, they can perform a range of tasks in unseen environments, generalize across scenes, viewpoints, objects, and language instructions [20, 43, 25, 50, 42, 31, 44, 15]. Though impressively general, these generalist policies often need to be adapted to perform effectively on downstream tasks or a new robot system, which is most commonly achieved by finetuning them on a curated dataset of demonstrations for the target task. While prior work has shown that such finetuning can lead to robust policies with tens or hundreds of hours of finetuning data [7, 20, 8, 9], collecting such amounts of robot demonstration data is challenging. As a result, in practice often less than 100 demonstrations or a few hours of robot data are used for finetuning [25, 51, 26]. Existing approaches for robot policy finetuning struggle to preserve the generality of the pre-trained model in such low-data regimes, and fail to robustly generalize far beyond the exact viewpoints, objects, and scenarios seen in the finetuning data [15, 62, 69, 57, 61, 71, 23]. To expand the usability of generalist policies, we need robust finetuning approaches that better preserve the generality of pre-trained robot policies and allow us to generalize to a broader set of scenarios on the target task.
In this work, we introduce RETAIN (Robust finE-tuning wiTh pArameter mergINg), a surprisingly simple approach for robust robot policy finetuning. We observe that, by simply interpolating the weights of the pre-trained generalist policy before and after finetuning on the target task (see Fig.˜1), we can obtain checkpoints that match the performance of the finetuned policy on scenarios present in the finetuning data, while generalizing significantly better to unseen variations of the target task, such as unseen object instances, positions, or viewpoints. Additionally, we observe that RETAIN preserves the generalist capabilities of the pre-trained policy also on tasks other than the target task, allowing us to use RETAIN in a continual learning setup by sequentially merging new skills into pre-trained generalist policies (in a literal sense). We demonstrate the effectiveness of RETAIN for robust policy finetuning and sequential skill acquisition across a range of real-world and simulated finetuning tasks, achieving state-of-the-art finetuning performance. While previous work has investigated interpolating model weights of pre-trained and fine-tuned models for vision and language [60, 59, 19], this is to our knowledge the first work to investigate and analyze parameter merging for robot policies and use it to enable continual acquisition of new robotic skills.
@@ 2 Related Work
Fueled by large-scale human teleoperated robot datasets [10, 52, 24, 46], generalist robot policies have recently been shown to solve a wide range of tasks across diverse scenes [7, 50, 42, 25, 9, 31, 1, 48, 44]. Yet, even state-of-the-art generalist policies typically need to be adapted for any given target task to achieve high performance [7]. Thus, a number of approaches have been proposed for training generalist policies on a new target task: from simply finetuning them on a dataset of target task demonstrations [7, 25, 50], or mixing the outputs of the pre-trained and fine-tuned policies [4], to alternative approaches like online and offline reinforcement learning [36, 18, 68, 17, 70, 63], retrieval-based adaptation [33, 12] or in-context improvement [13, 49]. In this work, we focus on the most common setting, in which policies are finetuned on a target task using a small dataset of demonstrations. Various finetuning approaches have been proposed in the literature, from simply adapting the full network on the target dataset [25, 7], to mixing target and pre-training data [8, 14, 11], or freezing parts of the network during finetuning [32, 29, 47, 64]. While such approaches may prove effective for learning robust target task policies in “large-data” finetuning regimes with tens to hundreds of hours of finetuning data [6, 20, 8, 9], they often struggle to retain the generality of the pre-trained policy in more common, accessible settings with 100 or less target task demonstrations. In such scenarios, finetuned policies often struggle to generalize meaningfully beyond the conditions seen in the finetuning dataset [69], even if the base policy had broad generalization capabilities. In this work, we propose a simple alternative for robust policy finetuning in low-data regimes. Instead of directly using the finetuned policy, we observe that merging the pretrained and finetuned policy checkpoints in weight space leads to significantly improved generalization on target tasks at no additional training or inference cost.
@@
Our approach is inspired by work on model weight merging in computer vision and natural language processing domains [55, 65, 66, 40, 35, 21, 37, 67, 22]. These works demonstrate that interpolating between the weights of multiple finetuned models, or between pre-trained and finetuned models, can combine their capabilities or make them more robust to distribution shifts [60, 59, 19, 41]. To our knowledge, our work is the first to demonstrate the effectiveness of model merging in the context of generalist robot policies, and combining it with co-training to further improve upon vanilla model merging. Additionally, we analyze the importance of different parameter groups in vision-language-action (VLA) policies and find it often sufficient to only merge parameters from the language model.
@@
The focus of our work is on improving generalization of finetuned policies on a target task. However, in addition, we find that our model merging approach is also effective at retaining the generalist policy’s performance on tasks from the pre-training distribution. As such, we demonstrate that it can be used to sequentially merge multiple skills into a single pre-trained policy checkpoint while retaining generality. This setting is typically referred to as continual learning and there is a large body of literature, both outside [27, 45, 34, 56] and within robotics [28, 3, 53, 38, 30, 58]. Our work differs from this line of research in that we aim to inherit and pass on the generalization ability of a pre-trained model to learn new tasks robustly, whereas continual learning methods generally focus on not forgetting old skills seen during the agent’s lifetime.
@@ 3 Problem Setting
The goal of our work is to develop an approach for robust policy finetuning, in which a generalist policy is finetuned to a new target task and generalizes to unseen variations of that target task, like new object instances, viewpoints, scenes, or lighting conditions, while also preserving its generalist abilities. Formally, let $\mathcal{M}$ denote an environment, $\mathcal{S}$ denote observations (e.g., images, proprioception), $\mathcal{A}$ actions, and $\mathcal{T}$ task specifications (e.g., language prompts). A policy $\pi_{\theta}(a_{t}\mid s_{t},T)$ maps state $s_{t}\in\mathcal{S}$ and task $T\in\mathcal{T}$ to a distribution over actions $a_{t}\in\mathcal{A}$. We assume access to a pretrained generalist policy $\pi_{\theta_{\mathrm{pre}}}$, trained on a diverse set of tasks and environments, and denote its training data as $\mathfrak{D}_{\mathrm{pre}}$. For a new target task $T_{\eta}$ (e.g., “wipe the whiteboard”), we assume access to a demonstration dataset $\mathfrak{D}_{\eta}\;=\;\big\{\,(s_{t}^{(i)},a_{t}^{(i)},T_{\eta})\,\big\}$. In general, we assume that $\mathfrak{D}_{\eta}$ is collected in a single (or small number) of environments $\mathcal{M}_{\eta}$, and $|\mathfrak{D}_{\eta}|\ll|\mathfrak{D}_{\mathrm{pre}}|$.
@@
We consider two primary finetuning settings: Task-finetuning (task-FT), in which we train exclusively on the target dataset $\mathfrak{D}_{\eta}$ (e.g., because the pre-training dataset is proprietary); and co-finetuning (co-FT), in which we finetune on a mix of $\mathfrak{D}_{\eta}$ and $\mathfrak{D}_{\mathrm{pre}}$ to help preserve pre-training capabilities (e.g., in case of open-source pre-training datasets).
@@
    Generalist tasks: measures policy performance on tasks other than the target task, but for which we would expect the generalist policy $\pi_{\theta_{\mathrm{pre}}}$ to perform reasonably. This measures how well the finetuned policy retains generalist capabilities from the pre-trained model.
@@ 4 Challenges of Finetuning Generalist Robot Policies
Figure 3: The standard approach for policy finetuning often overfits. As the policy is trained for more gradient steps, it perform worse on tasks other than the new target task (“GENERALIST”) and even start to degrade on scenarios seen in the finetuning data (“ID”). Most importantly, it is not able to transfer the generality of a base policy to do well under variations of the target task (new object positions, instances, viewpoints; “OOD”).
To understand the challenges of robust finetuning of generalist robot policies, we start by evaluating the standard finetuning approach. We evaluate finetuning the full model (Task-FT) in Libero [30], a multi-task robotic manipulation simulator containing 130 total tasks. Concretely, given a state-of-the-art vision-language-action policy [7], pre-trained on demonstration data from 117 tasks from the LIBERO-{90, goal, spatial, object} suites, we finetune it to a new Libero target task mugs-on-plates. We then measure performance in the three scenarios introduced in Section˜3: ID, OOD, and Generalist tasks. For OOD evaluations on the target tasks, we alter object positions, add new distractors, and change backgrounds. Generalist evaluations are performed over $20$ tasks from the pretraining dataset.
Fig.˜3 shows the performance of Task-FT on three types of evaluations. We find a clear tradeoff between generalist and ID target task performance: though the model gets better in ID target task after fine-tuning, it increasingly loses its generalist capabilities. In addition, when the model is finetuned for too long, it even starts losing performance for ID tasks. Both of these phenomena are likely because the model has severely overfitted to the small demonstration dataset, and is unable to do other tasks or recover from small mistakes unseen in the dataset. More importantly, one would hope that the finetuned model can generalize to small variations of the target task not seen in the finetuning dataset, since the pretrained model contains knowledge for such generalization. However, all tested checkpoints above show a large gap between ID and OOD performance, showing that the model is not able to complete the task when there are small variations to the target task. Section˜A.3 shows that even with more careful tuning of learning rate and number of gradient steps, such overfitting still exists.
@@ 5 RETAIN: Robust Policy Finetuning via Model Merging
The previous section illustrates that standard finetuning approaches quickly cause the model to forget its generalist capabilities, and fail to transfer the pre-trained policy’s robustness to the target task. To address these two issues, we propose RETAIN (Robust finE-tuning wiTh pArameter mergINg), a simple approach for robust finetuning of robot policies. Given pre-trained policy weights $\theta_{\mathrm{pre}}$ and finetuned policy weights $\theta_{\mathrm{ft}}$, we propose linearly interpolating $\theta_{\mathrm{pre}}$ and $\theta_{\mathrm{ft}}$ in weight space to obtain a final policy checkpoint, $\tilde{\theta}$. That is, RETAIN produces a final policy $\pi_{\tilde{\theta}}$ by setting:
@@
for $\alpha$ a tunable merging weight. Though surprisingly simple, as we will see, this weight space “merging” of pre-trained and finetuned checkpoints leads to significantly improved OOD performance on the target task, while retaining generalist policy capabilities (see Section˜6.4). While weight merging itself already improves the policy’s ability to retain and pass on generalist abilities, in the following we introduce two further improvements: utilizing the pretraining data $\mathfrak{D}_{\mathrm{pre}}$ (in settings where it is available) to augment our task data $\mathfrak{D}_{\eta}$ during finetuning, and merging $\theta_{\mathrm{pre}}$ and $\theta_{\mathrm{ft}}$ in a modality-specific manner. We will introduce these two methods below, and show how RETAIN can also enable continual adaptation to new tasks.
@@ 5.1 Co-Finetuning
In Eq.˜2, the finetuned policy weight $\theta_{\mathrm{ft}}$ can either be optimized via task-finetuning or co-finetuning, as described in Section˜3. In situations where the pre-training dataset, or a subset of it, is available, we can finetune the policy weight on a mix of $\mathfrak{D}_{\text{pre}}$ and $\mathfrak{D}_{\eta}$. Such co-finetuning usually leads to better retention of generalist abilities after finetuning [8, 14, 11]. While co-finetuning helps not forget prior knowledge, it is not effective at utilizing the prior knowledge to generalize in a new target task (see Fig.˜3). Instead, we find that we can use model merging together with co-finetuning, which we will refer to as RETAIN-co-FT, to enable greater generalization on the target task and better preserve generalist knowledge than model merging with task fine-tuning, which we call RETAIN-task-FT (see Section˜6).

### v2

@@ Abstract
Generalist robot policies, trained on large and diverse datasets, have demonstrated the ability to generalize across a wide spectrum of behaviors, enabling a single policy to act in varied real-world environments. However, they still fall short on new tasks not covered in the training data. When finetuned on limited demonstrations of a new task, these policies often overfit to the specific demonstrations—not only losing their prior abilities to solve a wide variety of generalist tasks but also failing to generalize within the new task itself. In this work, we aim to develop a method that preserves the generalization capabilities of the generalist policy during finetuning, allowing a single policy to robustly incorporate a new skill into its repertoire. Our goal is a single policy that both learns to generalize to variations of the new task and retains the broad competencies gained from pretraining. We show that this can be achieved through a simple yet effective strategy: interpolating the weights of a finetuned model with that of the pretrained model. We show, across extensive simulated and real-world experiments, that such model merging produces a single model that inherits the generalist abilities of the base model and learns to solve the new task robustly, outperforming both the pretrained and finetuned model on out-of-distribution variations of the new task. Moreover, we show that model merging performance scales with the amount of pretraining data, and enables continual acquisition of new skills in a lifelong learning setting, without sacrificing previously learned generalist abilities.
@@ 1 Introduction
Generalist robot policies trained on large corpora of data have recently shown impressive generalization abilities: out of the box, they can perform a range of tasks in unseen environments, generalize across scenes, viewpoints, objects, and language instructions [21, 45, 26, 52, 44, 32, 46, 16, 1]. Though impressively general, these generalist policies often need to be adapted to perform effectively on downstream tasks or a new robot system, which is most commonly achieved by finetuning them on a curated dataset of demonstrations for the target task. While prior work has shown that such finetuning can lead to robust policies with tens or hundreds of hours of finetuning data [8, 21, 9, 10], collecting such amounts of robot demonstration data is challenging. As a result, in practice often less than 100 demonstrations or a few hours of robot data are used for finetuning [26, 53, 27]. Existing approaches for robot policy finetuning struggle to preserve the generality of the pretrained model in such low-data regimes, and fail to robustly generalize far beyond the exact viewpoints, objects, and scenarios seen in the finetuning data [16, 64, 71, 59, 63, 73, 24]. To expand the usability of generalist policies, we need robust finetuning approaches that better preserve the generality of pretrained robot policies and allow us to generalize to a broader set of scenarios on the target task.
In this work, we introduce RETAIN (Robust finE-tuning wiTh pArameter mergINg), a surprisingly simple approach for robust robot policy finetuning. By simply interpolating the weights of the pretrained generalist policy before and after finetuning on the target task (see Fig.˜1), we obtain checkpoints that match the performance of the finetuned policy on scenarios present in the finetuning data, while generalizing significantly better to unseen variations of the target task, such as unseen object instances, positions, or viewpoints. Additionally, we observe that RETAIN preserves the generalist capabilities of the pretrained policy also on tasks other than the target task, allowing us to use RETAIN in a continual learning setup by sequentially merging new skills into pretrained generalist policies (in a literal sense). We demonstrate the effectiveness of RETAIN for robust policy finetuning and sequential skill acquisition across a range of real-world and simulated finetuning tasks, achieving state-of-the-art finetuning performance. We also show that RETAIN gets even more effective when the pretrained policy is trained on more data. While previous work has investigated interpolating model weights of pretrained and fine-tuned models for vision and language [62, 61, 20], this is to our knowledge the first work to investigate and analyze parameter merging for robot policies and use it to enable continual acquisition of new robotic skills.
@@ 2 Related Work
Fueled by large-scale human teleoperated robot datasets [11, 54, 25, 48], generalist robot policies have recently solved a wide range of tasks across diverse scenes [8, 52, 44, 26, 10, 32, 2, 50, 46]. Yet, even state-of-the-art generalist policies typically need to be adapted for any given target task to achieve high performance [8]. Thus, a number of approaches have been proposed for training generalist policies on a new target task: from simply finetuning them on a dataset of target task demonstrations [8, 26, 52], or mixing the outputs of the pretrained and fine-tuned policies [5], to alternative approaches like online and offline reinforcement learning [38, 19, 70, 18, 72, 65], retrieval-based adaptation [34, 13] or in-context improvement [14, 51]. In this work, we focus on the most common setting, in which policies are finetuned on a target task using a small dataset of demonstrations. Various finetuning approaches have been proposed in the literature, from simply adapting the full network on the target dataset [26, 8], to mixing target and pretraining data [9, 15, 12], or freezing parts of the network during finetuning [33, 30, 49, 66]. While such approaches may prove effective for learning robust target task policies in “large-data” finetuning regimes with tens to hundreds of hours of finetuning data [7, 21, 9, 10], they often struggle to retain the generality of the pretrained policy in more common, accessible settings with 100 or less target task demonstrations. In such scenarios, finetuned policies often struggle to generalize meaningfully beyond the conditions seen in the finetuning dataset [71], even if the base policy had broad generalization capabilities. In this work, we propose a simple alternative for robust policy finetuning in low-data regimes. Instead of directly using the finetuned policy, we observe that merging the pretrained and finetuned policy checkpoints in weight space leads to significantly improved generalization on target tasks at no additional training or inference cost.
@@
Our approach is inspired by work on model weight merging in vision and language domains [57, 67, 68, 42, 36, 22, 39, 69, 23]. These works demonstrate that interpolating between the weights of multiple finetuned models, or between pretrained and finetuned models, can combine their capabilities or make them more robust to distribution shifts [62, 61, 20, 43, 37]. To our knowledge, our work is the first to demonstrate the effectiveness of model merging in the context of generalist robot policies, and combining it with co-training to further improve upon vanilla model merging. Additionally, we analyze the importance of the generality of the base model, as well as the importance of different parameter groups in vision-language-action (VLA) policies and find it often sufficient to only merge parameters from the language model backbone.
@@
The focus of our work is on improving generalization of finetuned policies on a target task. However, in addition, we find that our model merging approach is also effective at retaining the generalist policy’s performance on tasks from the pretraining distribution. As such, we demonstrate that it can be used to sequentially merge multiple skills into a single pretrained policy checkpoint while retaining generality. This setting is typically referred to as continual learning and there is a large body of literature, both outside [28, 47, 35, 58] and within robotics [29, 4, 55, 40, 31, 60]. Our work differs from this line of research in that we aim to inherit and pass on the generalization ability of a pretrained model to learn new tasks robustly, whereas continual learning methods generally focus on not forgetting old skills seen during the agent’s lifetime.
@@ 3 Problem Setting
The goal of our work is to develop an approach for robust policy finetuning, in which a generalist policy is finetuned to a new target task and generalizes to unseen variations of that target task, like new object instances, viewpoints, scenes, or lighting conditions, while also preserving its generalist abilities. Formally, let $\mathcal{M}$ denote an environment, $\mathcal{S}$ denote observations (e.g., images, proprioception), $\mathcal{A}$ actions, and $\mathcal{T}$ task specifications (e.g., language prompts). A policy $\pi_{\theta}(a_{t}\mid s_{t},T)$ maps state $s_{t}\in\mathcal{S}$ and task $T\in\mathcal{T}$ to a distribution over actions $a_{t}\in\mathcal{A}$. We assume access to a pretrained generalist policy $\pi_{\theta_{\mathrm{pre}}}$, trained on a diverse set of tasks and environments, and denote its training data as $\mathfrak{D}_{\mathrm{pre}}$. For a new target task $T_{\eta}$ (e.g., “wipe the whiteboard”), we assume access to a demonstration dataset $\mathfrak{D}_{\eta}\;=\;\big\{\,(s_{t}^{(i)},a_{t}^{(i)},T_{\eta})\,\big\}$. In general, we assume that $\mathfrak{D}_{\eta}$ is collected in a single (or small number of) environment $\mathcal{M}_{\eta}$, and $|\mathfrak{D}_{\eta}|\ll|\mathfrak{D}_{\mathrm{pre}}|$.
@@
We consider two main finetuning settings: Task-finetuning (task-FT), in which we train exclusively on the target dataset $\mathfrak{D}_{\eta}$ (e.g., because pretraining data is proprietary); and co-finetuning (co-FT), where we finetune on a mix of $\mathfrak{D}_{\eta}$ and $\mathfrak{D}_{\mathrm{pre}}$ to help preserve pretraining capabilities (e.g., in case of open-source pretraining datasets).
Figure 3: Example filmstrips of the in-distribution (ID) tasks and out-of-distribution (OOD) tasks from DROID robot experiments in Section˜6.
@@
    Generalist tasks: measures policy performance on tasks other than the target task, but for which we would expect the generalist policy $\pi_{\theta_{\mathrm{pre}}}$ to perform reasonably. This measures how well the finetuned policy retains generalist capabilities from the pretrained model.
@@ 4 Challenges of Finetuning Generalist Robot Policies
To understand the challenges of robust finetuning of generalist robot policies, we start by evaluating the standard finetuning approach. We evaluate finetuning the full model (Task-FT) in LIBERO [31], a multi-task robotic manipulation simulator containing 130 total tasks. Concretely, given a state-of-the-art vision-language-action policy [8], pretrained on demonstration data from 117 tasks from the LIBERO-{90, goal, spatial, object} suites, we finetune it to a new LIBERO target task. We consider three target tasks from LIBERO-10: mugs-on-plates, pot-on-stove, and items-into-basket. We then measure performance in the three scenarios introduced in Section˜3: ID, OOD, and Generalist. For OOD evaluations, we alter object positions, add new distractors, and change backgrounds. Generalist evaluations are performed over $20$ tasks from the pretraining dataset. More details about the OOD evaluations in Section˜6.1 and Section˜A.6.
Figure 4: The standard approach for policy finetuning often overfits. As the policy is trained for more gradient steps, it performs worse on tasks other than the new target task (“GENERALIST”) and may even start to degrade on scenarios seen in the finetuning data (“ID”). Most importantly, it is not able to transfer the generality of a base policy to do well under variations of the target task (new object positions, instances, viewpoints; “OOD”).
Fig.˜4 shows the performance of Task-FT on three types of evaluations. We find a clear tradeoff between generalist and ID target task performance: though the model gets better in ID target task after fine-tuning, it increasingly loses its generalist capabilities. Additionally, when the model is finetuned for too long, it even starts losing performance for ID tasks. Both of these phenomena are likely because the model has severely overfitted to the small demonstration dataset, and is unable to do other tasks or recover from small mistakes unseen in the dataset. More importantly, one would hope that the finetuned model can generalize to small variations of the target task not seen in the finetuning dataset, since the pretrained model contains knowledge for such generalization. However, all tested checkpoints above show a large gap between ID and OOD performance, showing that the model is not able to complete the target task when there are small variations present. Section˜A.3 shows that even with careful tuning of learning rate and number of gradient steps, such overfitting still exists.
@@ 5 RETAIN: Robust Policy Finetuning via Model Merging
The previous section illustrates that standard finetuning approaches cause the model to quickly forget generalist capabilities, and fail to transfer the pretrained policy’s robustness to the target task. To address these issues, we propose RETAIN (Robust finE-tuning wiTh pArameter mergINg), a simple approach for robust finetuning of robot policies. Given pretrained policy weights $\theta_{\mathrm{pre}}$ and finetuned policy weights $\theta_{\mathrm{ft}}$, we propose linearly interpolating $\theta_{\mathrm{pre}}$ and $\theta_{\mathrm{ft}}$ to obtain a final policy checkpoint, $\tilde{\theta}$. So, RETAIN produces a final policy $\pi_{\tilde{\theta}}$ by setting:
@@
for $\alpha$, a tunable merging weight. Though surprisingly simple, as we will see, this weight space “merging” of pretrained and finetuned checkpoints leads to significantly improved OOD performance on the target task, while retaining generalist policy capabilities (see Section˜6.5). While weight merging itself already improves the policy’s ability to retain and pass on generalist abilities, in the following we introduce two further improvements: utilizing the pretraining data $\mathfrak{D}_{\mathrm{pre}}$ (in settings where it is available) to augment our task data $\mathfrak{D}_{\eta}$ during finetuning, and merging $\theta_{\mathrm{pre}}$ and $\theta_{\mathrm{ft}}$ in a modality-specific manner. We introduce these two methods below, and show how RETAIN can also enable continual adaptation to new tasks.
@@ 5.1 Co-Finetuning
In Eq.˜2, the finetuned policy weight $\theta_{\mathrm{ft}}$ can either be optimized via task-finetuning or co-finetuning, as described in Section˜3. In situations where the pretraining dataset, or a subset of it, is available, we can finetune the policy weight on a mix of $\mathfrak{D}_{\text{pre}}$ and $\mathfrak{D}_{\eta}$. Such co-finetuning usually leads to better retention of generalist abilities after finetuning [9, 15, 12]. We find that we can use model merging together with co-finetuning, which we will refer to as RETAIN-co-FT, to enable greater generalization on the target task and better preserve generalist knowledge than model merging with task fine-tuning, which we call RETAIN-task-FT (see Section˜6).